GLOBAL VIBRATION-ROTATION ANALYSIS IN ACETYLENE: ^{12}CH^{13}CH

S. ROBERT, \underline{M. HERMAN}, Service de Chimie quantique et Photophysique, CP160/09, Université Libre de Bruxelles (U.L.B.), Ave. F.-D. Roosevelt, 50, B-1050, Bruxelles, Belgium; A. FAYT, Laboratoire de Spectroscopie Moléculaire, Université Catholique de Louvain, Chemin du Cyclotron 2, B-1348 Louvain-La-Neuve, Belgium; G. DI LONARDO, L. FUSINA, and F. TAMASSIA, Dipartimento di Chimica Fisica e Inorganica, Università di Bologna, Viale Risorgimento, 4, I-40136, Bologna, Italy.

All known vibration-rotation absorption lines of ^{12}CH^{13}CH accessing levels up to 6750 cm^{−1} were gathered from the literature. They were fitted simultaneously to J-dependent Hamiltonian matrices exploiting the well known vibrational polyad or cluster block-diagonalization, in terms of the pseudo quantum numbers N_{s} = v_{1}+v_{2}+v_{3} and N_{r} = 5v_{1}+3v_{2}+5v_{3}+v_{4}+v_{5}, and accounting also for l-parity and e/f symmetry properties. The anharmonic interaction coupling terms known to occur from a pure vibrational fit in this acetylene isotopologue were included in the model. The results will be presented and discussed.